how do i use the zero-factor property when solving a quadratic equation. can they both sides be zero?
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

for example,
(x + 1)*(x + 2) = 0
For for product to be zero, either x + 1 = 0, or x + 2 = 0.
If both terms are not zero, then the product would not be zero.
-------------
This does not apply to other numbers.
(x+1)*(x+2) = 4 does NOT mean than one of the terms = 4, or that they both equal 2, or any other limitation.
It's the "zero product", the product must equal zero.

Answer by stanbon(75887)

how do i use the zero-factor property when solving a quadratic equation. can they both sides be zero?
----
I'm not sure what you mean be "both sides be zero".
If both sides are zero you do not have a quadratic equation.
----
Maybe you mean, "Can both of the solutions be zero?".
---
The answer to that is yes.
The quadratic y = x^2 has 2 Real zeros.
They are both zero.
==========================
Cheers,
Stan H.